# Learned harmonic mean estimation of the Bayesian evidence with normalizing flows

### v2

@@ Abstract
[…]
@@ 2.1 Bayesian model comparison
Bayes’ theorem is a fundamental principle in Bayesian statistics that allows us to update our beliefs about models in light of observed data. Consider observed data ${y}$ described through a model $M$ parametrised by $\theta$. Bayes’ theorem gives us the posterior ${p}(\theta{\,|\,}{y},M)$, the probability density of a model’s parameter $\theta$ given observed data ${y}$ and model $M$. It is expressed in terms of the prior probability density of the model, the likelihood of the data under that model, and the Bayesian evidence for the data:
@@
The likelihood ${p}({y}{\,|\,}\theta,M)=\mathcal{L}(\theta)$ expresses how probable the observed data $y$ is for different values of the parameter $\theta$. The prior ${p}(\theta{\,|\,}M)=\pi(\theta)$ quantifies pre-existing knowledge or assumptions about $\theta$. The Bayesian evidence, also called the marginal likelihood, ${p}({y}{\,|\,}M)=z$ is a normalizing factor for the posterior distribution.
@@ 3.2.1 Training the flow
$=$ $-\mathbb{E}_{p(\theta)}\biggl{[}\log\left(\,p_{\text{NF}}(\theta,\beta)\,\right)\biggr{]}+\text{const.}$. (17)
@@
$L(\beta)\approx-\frac{1}{N}\sum_{i=1}^{N}\log\left(\,p_{\text{NF}}(\theta_{i},\beta)\,\right)+\text{const.}$. (18)
@@ 3.2.4 Evidence error estimate
$\hat{\sigma}/\hat{\rho}=\text{exp}\bigl{(}\log(\hat{\sigma})-\log(\hat{\rho})\bigr{)}.$ (23)
This way we can avoid computing $\hat{\rho}\pm\hat{\sigma}$ directly. We only compute $\log(\hat{\sigma})-\log(\hat{\rho})$, which we expect to be much smaller and less susceptible to overflow. When quoting the result with log-space errors we use the notation $\log(\hat{\rho})^{\hat{\zeta}_{+}}_{\hat{\zeta}_{-}}$. The log evidence errors can be straightforwardly obtained by swapping the negative and positive errors of the reciprocal log evidence.
@@ 4.1 Architectures, sampling and training
In the experiments where a real NVP flow is used, translation networks of the affine coupling layers are given by two-layer dense neural networks with a leaky ReLU activation function in between. For the scaling layers this is scaled by another dense layer with a softplus activation. We permute the inputs between coupling layers to ensure the flow transforms all elements. In the low dimensional benchmark experiments we consider a real NVP flow, unless otherwise stated, with six coupling layers, where typically only the first two include scaling. When we use a rational quadratic spline flow, it has a range $-10$ to $10$ (outside of this range it defaults to a linear transformation). The conditioner for the spline hyperparameters is a multi-layer perceptron with a hyperbolic tangent activation. We use a Gaussian base distribution with zero mean and an identity covariance matrix for all flows.
@@ 4.2.3 Logistic regression models: Pima Indian example
Model $\text{Model }M_{1}:$ covariates = {NP, PGC, BMI, DP};
Model $\text{Model }M_{2}:$ covariates = {NP, PGC, BMI, DP, AGE},
@@
Figure 5 shows the corner plots for this example for both models. The training samples from the posterior are shown in red and from the normalizing flow at temperature $T=0.9$ in blue. Once again, we see that the concentrated flow is contained within the posterior as expected. The log evidence found for Model 1 and 2 is $-257.230^{0.003}_{-0.003}$ and $-259.857^{0.002}_{-0.002}$ respectively, resulting in the estimate $\log\text{BF}_{12}=2.627^{0.004}_{-0.004}$, indicating a slight preference for Model 1. The Bayes factor value is in close agreement with the benchmark, whereas the original harmonic mean estimator was not accurate (Friel & Wyse, 2012).
@@ 4.2.4 Non-nested linear regression models: Radiata pine example
Model $\text{Model }M_{1}$ ::: $y_{i}$ $=\alpha+\beta(x_{i}-\bar{x})+\epsilon_{i},$ $\epsilon_{i}$ $\sim\text{N}(0,\tau^{-1});$ (31)
Model $\text{Model }M_{2}$ ::: $y_{i}$ $=\gamma+\delta(z_{i}-\bar{z})+\eta_{i},$ $\eta_{i}$ $\sim\text{N}(0,\lambda^{-1}),$ (32)
@@
Figure 6 shows a corner plot of the training samples from the posterior (red) and from the normalizing flow (blue) at temperature $T=0.9$ for both models. Again, it can be seen that the concentrated learned target is contained within the posterior. The log evidence found for Model 1 and 2 is $-310.1284^{0.0007}_{-0.0007}$ and $-301.7044^{0.0008}_{-0.0008}$ respectively, resulting in the estimate $\log\text{BF}_{12}=8.424^{0.001}_{-0.001}$. The analytic values of the log evidence are $-310.1283$ and $-301.7046$ for Models 1 and 2 respectively, resulting in the estimate $\log\text{BF}_{12}=8.424$. The value obtained using our estimator is in close agreement with the ground truth. The learned harmonic mean gives an accurate estimate of the evidence, whereas the original harmonic mean estimator fails catastrophically for this example (Friel & Wyse, 2012).
@@ 4.4 Robustness of the temperature parameter
Learned harmonic mean $-65.262_{-0.011}^{+0.011}$ $-67.407_{-0.009}^{0.009}$ $2.145_{-0.014}^{0.014}$ 16 hours (sampling) $+$ 16 minutes (evidence)
Nested sampling $-65.21\pm 0.32$ $-67.44\pm 0.32$ $2.23\pm 0.45$ 94 hours (sampling and evidence)
Naïve flow estimator $-64.9\pm 0.8$ $-67.0\pm 1.1$ $2.1\pm 1.4$ Similar to learned harmonic mean
@@ 4.5 Practical cosmological example: DES Y1 analysis
(a) $\Lambda$CDM
(b) $w$CDM
Figure 10: Corner plot of the sampled posterior (solid red) and rational quadratic spline flow trained on the posterior samples with temperature $T=0.8$ (dashed blue) for the DES Y1 analysis example for (a) $\Lambda$CDM and (b) $w$CDM cosmological models. The internal importance target distribution given by the concentrated flow is contained within the posterior and has thinner tails, as required for the learned harmonic mean estimator, even in this higher dimensional case.
@@
We follow the reference approach described by Campagne et al. (2023), who extract and compress a subset of the DES Year 1 lensing and clustering data and set up a forward model following the DES Y1 Pipeline (Abbott et al., 2018a). We use the DES Y1 redshift distributions [^4] http://desdr-server.ncsa.illinois.edu/despublic/y1a1_files/chains/2pt_NG_mcal_1110.fits and simulate a 3x2pt data vector for a fixed cosmology. We use this as our mock data vector and run an inference pipeline to obtain posterior contours and evidence estimates. We refer the reader to Campagne et al. (2023) for the priors and all other details. To sample from the posterior we use the Cobaya package (Torrado & Lewis, 2019) with the Metropolis-Hastings algorithm. We then apply harmonic to these samples to evaluate the Bayesian evidence. For comparison, we also sample using the PolyChord nested sampler (Handley et al., 2015a, b) in Cobaya, which provides a benchmark Bayesian evidence estimate. We perform the analysis twice, assuming either a $\Lambda$CDM or $w$CDM cosmological model, with the dark energy equation of state parameter $w$ fixed to $w=-1$ or free to vary, respectively.
The $\Lambda$CDM and $w$CDM models have $20$ and $21$ parameters respectively. With the Metropolis-Hastings sampler, we run $64$ chains per model, obtaining an average of approximately $5,800$ and $6,500$ samples per chain for $\Lambda$CDM and $w$CDM respectively. We discard $500$ samples for burn-in in both cases. We train a rational quadratic flow consisting of $3$ layers, with $128$ spline bins, applying standardization, on half of the chains, and use the other half for inference.
@@
Note that the values computed by the learned harmonic mean and nested sampling are in agreement, showing a slight preference for $\Lambda$CDM, matching the configuration of our simulated setup. The values computed by the naïve estimator are in approximate agreement but exhibit an error two orders or magnitude larger than the error of the learned harmonic mean (in higher dimensional examples to be reported in an ongoing work we observe the naïve estimator failing much more catastrophically).
In terms of computational speed (summarized in Table 1 but reported in great detail here), sampling with Cobaya using the Metropolis-Hastings algorithm takes approximately $8$ hours for $\Lambda$CDM and $w$CDM each on $64$ CPU cores. The compute time added by harmonic is around $5$ minutes on $1$ GPU for training and $3$ minutes on $128$ CPU cores to estimate the evidence for each model. Using PolyChord takes approximately $47$ hours for $\Lambda$CDM and $w$CDM each, on the same $64$ CPU cores used for the Metropolis-Hastings sampling. The Metropolis-Hastings algorithm in this case is much quicker due to the use of a proposal covariance matrix based on a Planck cosmology (Campagne et al., 2023). Thanks to the flexibility of the learned harmonic estimator, we can leverage this advantage and choose Metropolis-Hastings over nested sampling, while still being able to estimate the Bayesian evidence and perform model comparison. Even in this higher dimensional setting, the learned harmonic mean only adds a few minutes of compute time on top of the sampler. This demonstrates the potential scalability of the method and its potential for computing the evidence from existing saved down MCMC chains.

### v3

@@ Abstract
We present the learned harmonic mean estimator with normalizing flows – a robust, scalable and flexible estimator of the Bayesian evidence for model comparison. Since the estimator is agnostic to sampling strategy and simply requires posterior samples, it can be applied to compute the evidence using any Markov chain Monte Carlo (MCMC) sampling technique, including saved down MCMC chains, or any variational inference approach. The learned harmonic mean estimator was recently introduced, where machine learning techniques were developed to learn a suitable internal importance sampling target distribution to solve the issue of exploding variance of the original harmonic mean estimator. In this article we present the use of normalizing flows as the internal machine learning technique within the learned harmonic mean estimator. Normalizing flows can be elegantly coupled with the learned harmonic mean to provide an approach that is more robust, flexible and scalable than the machine learning models considered previously. We perform a series of numerical experiments, applying our method to benchmark problems and to a cosmological example in up to $21$ dimensions. We find the learned harmonic mean estimator is in agreement with ground truth values and nested sampling estimates. The open-source harmonic Python package implementing the learned harmonic mean, now with normalizing flows included, is publicly available.
@@ 2.1 Bayesian model comparison
Bayes’ theorem is a fundamental principle in Bayesian statistics that allows us to update our beliefs about models in light of observed data. Consider observed data y{y} described through a model $M$ parametrised by $\theta$. Bayes’ theorem gives us the posterior ${p}(\theta{\,|\,}{y},M)$, the probability density of a model’s parameter $\theta$ given observed data y{y} and model $M$. It is expressed in terms of the prior probability density of the model, the likelihood of the data under that model, and the Bayesian evidence for the data:
@@
The likelihood ${p}({y}{\,|\,}\theta,M)=\mathcal{L}(\theta)$ expresses how probable the observed data $y$ is for different values of the parameter $\theta$. The prior ${p}(\theta{\,|\,}M)=\pi(\theta)$ quantifies pre-existing knowledge or assumptions about $\theta$. The Bayesian evidence, also called the marginal likelihood, p(y|M)=z{p}({y}{\,|\,}M)=z is a normalizing factor for the posterior distribution.
@@ 3.2.1 Training the flow
$=$ $-\mathbb{E}_{p(\theta)}\biggl[\log\left(\,p_{\text{NF}}(\theta,\beta)\,\right)\biggr]+\text{const.}$ (17)
@@
$L(\beta)\approx-\frac{1}{N}\sum_{i=1}^{N}\log\left(\,p_{\text{NF}}(\theta_{i},\beta)\,\right)+\text{const.}$ (18)
@@ 3.2.4 Evidence error estimate
[…]
@@ 4.1 Architectures, sampling and training
In the experiments where a real NVP flow is used, translation networks of the affine coupling layers are given by two-layer dense neural networks with a leaky ReLU activation function in between. For the scaling layers this is scaled by another dense layer with a softplus activation. We permute the inputs between coupling layers to ensure the flow transforms all elements. In the low dimensional benchmark experiments we consider a real NVP flow, unless otherwise stated, with six coupling layers, where typically only the first two include scaling. When we use a rational quadratic spline flow, it has a range 10-10 to $10$ (outside of this range it defaults to a linear transformation). The conditioner for the spline hyperparameters is a multi-layer perceptron with a hyperbolic tangent activation. We use a Gaussian base distribution with zero mean and an identity covariance matrix for all flows.
@@ 4.2.3 Logistic regression models: Pima Indian example
$\text{Model }M_{1}:$ covariates = {NP, PGC, BMI, DP};
$\text{Model }M_{2}:$ covariates = {NP, PGC, BMI, DP, AGE},
@@
Figure 5 shows the corner plots for this example for both models. The training samples from the posterior are shown in red and from the normalizing flow at temperature $T=0.9$ in blue. Once again, we see that the concentrated flow is contained within the posterior as expected. The log evidence found for Model 1 and 2 is $257.230^{0.003}_{-0.003}$ and $259.857^{0.002}_{-0.002}$ respectively, resulting in the estimate $\log\text{BF}_{12}=2.627^{0.004}_{-0.004}$, indicating a slight preference for Model 1. The Bayes factor value is in close agreement with the benchmark, whereas the original harmonic mean estimator was not accurate (Friel & Wyse, 2012).
@@ 4.2.4 Non-nested linear regression models: Radiata pine example
$\text{Model }M_{1}$ $:$ $y_{i}$ $=\alpha+\beta(x_{i}-\bar{x})+\epsilon_{i},$ $\epsilon_{i}$ $\sim\text{N}(0,\tau^{-1});$ (31)
$\text{Model }M_{2}$ $:$ $y_{i}$ $=\gamma+\delta(z_{i}-\bar{z})+\eta_{i},$ $\eta_{i}$ $\sim\text{N}(0,\lambda^{-1}),$ (32)
@@
Figure 6 shows a corner plot of the training samples from the posterior (red) and from the normalizing flow (blue) at temperature $T=0.9$ for both models. Again, it can be seen that the concentrated learned target is contained within the posterior. The log evidence found for Model 1 and 2 is $310.1284^{0.0007}_{-0.0007}$ and $301.7044^{0.0008}_{-0.0008}$ respectively, resulting in the estimate $\log\text{BF}_{12}=8.424^{0.001}_{-0.001}$. The analytic values of the log evidence are 310.1283-310.1283 and 301.7046-301.7046 for Models 1 and 2 respectively, resulting in the estimate $\log\text{BF}_{12}=8.424$. The value obtained using our estimator is in close agreement with the ground truth. The learned harmonic mean gives an accurate estimate of the evidence, whereas the original harmonic mean estimator fails catastrophically for this example (Friel & Wyse, 2012).
@@ 4.4 Robustness of the temperature parameter
Learned harmonic mean $65.262_{-0.011}^{+0.011}$ $67.407_{-0.009}^{0.009}$ $2.145_{-0.014}^{0.014}$ 16 hours (sampling) ++ 16 minutes (evidence)
Nested sampling $65.21\pm 0.32$ $67.44\pm 0.32$ $2.23\pm 0.45$ 94 hours (sampling and evidence)
Naïve flow estimator $64.9\pm 0.8$ $67.0\pm 1.1$ $2.1\pm 1.4$ Similar to learned harmonic mean
@@ 4.5 Practical cosmological example: DES Y1 analysis
[…]
(b) wwCDM
[…]
@@
[…]
@@
[…]